Stocks A and B each have an expected return of 15%, a standard deviation of 20%, and a beta of 1.2. The returns on the two stocks have a correlation coefficient of -1.0. You have a portfolio that consists of 50% A and 50% B. Which of the following statements is CORRECT?

The portfolio's standard deviation is zero (i.e., a riskless portfolio).

The portfolio's beta is greater than 1.2.

The portfolio's standard deviation is greater than 20%.

The portfolio's expected return is less than 15%.

The portfolio's beta is less than 1.2.
Answer #1

А C D E F G H I J K L M N O P Q R Expected Return Standard Deviation А 15% 20% в 15% 20% Beta = 1.2 Correlation = -1.0 Expect
Answer #2

ANSWER :



Beta of portfolio = WA bA+ WB bB = 0.5*1.2 + 0.5*1.2 = 1.2 . 

So, option 2 and 5  are ruled out.


Expected return of portfolio = WA rA + WB rB = 0.5*15 + 0.5*15 = 15% . 

So, option 4 is ruled out.


SD of portfolio = sqrt ((WA sA)^2 + (WB sB)^2 + 2WA WB sA sB r)

= sqrt((0.5*0.2)^2 + (0.5*0.2)^2 + 2*0.5*0.5*0.2*0.2*(-1.0))

= 0

So, option 1 is correct and option 3 is ruled out.



Therefore, the option 1 : The portfolio’s SD is zero (risk-less)  is CORRECT.

(ANSWER


answered by: Tulsiram Garg